# EPRecon: An Efficient Framework for Real-Time Panoptic 3D Reconstruction from Monocular Video

### v1

@@ Abstract
Panoptic 3D reconstruction from a monocular video is a fundamental perceptual task in robotic scene understanding. However, existing efforts suffer from inefficiency in terms of inference speed and accuracy, limiting their practical applicability. We present EPRecon, an efficient real-time panoptic 3D reconstruction framework. Current volumetric-based reconstruction methods usually utilize multi-view depth map fusion to obtain scene depth priors, which is time-consuming and poses challenges to real-time scene reconstruction. To end this, we propose a lightweight module to directly estimate scene depth priors in a 3D volume for reconstruction quality improvement by generating occupancy probabilities of all voxels. In addition, to infer richer panoptic features from occupied voxels, EPRecon extracts panoptic features from both voxel features and corresponding image features, obtaining more detailed and comprehensive instance-level semantic information and achieving more accurate segmentation results. Experimental results on the ScanNetV2 dataset demonstrate the superiority of EPRecon over current state-of-the-art methods in terms of both panoptic 3D reconstruction quality and real-time inference. Code is available at https://github.com/zhen6618/EPRecon.
@@ I INTRODUCTION
Recently, visual-based panoptic 3D reconstruction [1] has gradually attracted attention due to its rich scene understanding information and wide application scenarios, such as robotic mapping and manipulation [2], and augmented reality (AR) [3]. Since camera poses can be accurately obtained by systems such as visual odometry [4], robot kinematics [5], or Simultaneous Localization and Mapping (SLAM) [6, 7], the main challenge in panoptic 3D reconstruction from a monocular video is how to obtain dense and high-quality scene surface depth from continuous RGB sequences and perform panoptic segmentation on the reconstructed scene.
@@
Methods that fuse multi-view depth maps obtained from similarity matching along epipolar lines, such as MVSNet series [8, 9, 10, 11, 12], suffer from depth estimation inconsistencies and redundant computations. To end this, volumetric-based methods [13, 14, 15, 16, 1] directly predict the explicit surface occupancy probabilities or implicit truncated signed distance function (TSDF) values to reconstruct scene surfaces in a 3D volume. To improve reconstruction quality, some works [16, 1, 17] first use off-the-shelf depth estimation models such as monocular depth estimation [18] for a preliminary estimate of scene depth priors, which can eliminate the majority of non-surface voxels (over 90%), and then perform depth-guided reconstruction. Although these methods effectively improve reconstruction performance, the scene depth estimation networks used in existing methods are time-consuming, posing challenges to real-time scene reconstruction.
@@
To simultaneously obtain panoptic segmentation information and depth information from image sequences, recent studies [1, 19] first predict occupancy probabilities of voxels, and then conduct 3D panoptic segmentation on predicted reconstructed voxels. Although these methods have shown certain efficacy, their panoptic features are only extracted from occupied voxel features or corresponding image features, resulting in limited instance-level semantic information. This is because voxel features need to generate rich geometric features for depth predictions and usually emphasize detailed semantic information. After multi-scale feature extraction and fusion, image features generally emphasize comprehensive semantic information.
@@
In this paper, we propose EPRecon, an efficient real-time panoptic 3D reconstruction framework. We first estimate scene depth priors in a 3D volume to eliminate most of non-surface voxels, and then perform depth-guided panoptic 3D reconstruction. Unlike previous works [16, 1, 17] which directly use time-consuming depth estimation models to obtain 2D depth priors and then perform depth fusion, EPRecon introduces a lightweight module to directly estimate scene depth priors in a 3D volume by generating occupancy probabilities of all voxels. Specifically, EPRecon back-projects multi-view image features into the volume and extracts feature similarities between views. These features are further aggregated and converted into occupancy probabilities of voxels. The module makes EPRecon focus more on voxel feature extraction in regions near scene surfaces and improves the entire panoptic reconstruction quality.
@@
  • We introduce a lightweight module to directly estimate scene depth priors in a 3D volume, which eliminates the majority of non-surface voxels and improves panoptic reconstruction quality.
  • We extract panoptic features from voxel features and corresponding image features, obtaining more detailed and comprehensive panoptic segmentation information and achieving more accurate segmentation results.
@@ II Related Work
3D reconstruction from RGB image sequences has a long research history, and many classic works have emerged, such as MVSNet [8] and ORB-SLAM3 [6]. Multi-view stereo reconstruction methods represented by MVSNet series [8, 9, 10, 11, 12] estimate depth map for each view and then perform depth map fusion. These methods greatly improve the accuracy of 3D reconstruction, however, multiple depth estimations from different views suffer from depth estimation inconsistencies and redundant computations, usually generating artifacts and repeated surfaces. Volumetric-based approaches [13, 14, 15, 21, 22, 23, 24, 25] directly predict surface occupancy probabilities or TSDF values in a 3D volume. An early work is Atlas [13], which directly regressed a TSDF volume from a set of posed RGB images. NeuralRecon [14] reconstructed local surfaces represented by sparse TSDF volumes for each video fragment sequentially, and used GRUs to fuse TSDF volumes. To achieve more detailed reconstruction, recent studies [26, 27, 1, 16, 17] first estimate scene depth priors and then perform depth-guided scene reconstruction. For example, DG-Recon [26] utilized multi-view depth priors to eliminate many irrelevant non-surface voxels and guide surface reconstruction. FineRecon [27] combined depth priors and global image features for surface occupancy predictions. PanoRecon [1] performed panoptic reconstruction with the guidance of MVS depth estimation. Although a preliminary estimate of the scene depth can recover more accurate surfaces, the depth prior prediction modules in existing methods are time-consuming, posing challenges to real-time panoptic reconstruction. To end this, we propose a lightweight module to directly estimate scene depth priors in the 3D volume.
@@ IV-C Ablation Studies
In Table V, we conduct ablation studies to verify the effectiveness of the components in EPRecon. The qualitative comparison and analysis are shown in Fig. 3. The proposed depth prior estimation module effectively improves the quality of 3D geometric reconstruction and 3D panoptic segmentation. Since we directly estimate scene depth priors in a 3D volume, our proposed module is more lightweight (13.4$\times$ faster) than current depth prior estimation methods based on depth map fusion (see Table IV). In addition, we compare with extracting panoptic features only from voxel features or corresponding image features. Our method captures both types of features, having more detailed and comprehensive instance-level semantic information, which leads to higher 3D panoptic segmentation performance.
TABLE IV: Runtime analysis of panoptic 3D reconstruction
@@
EPRecon 32 ms 287 ms 31.36 KFPS

### v2

@@ Abstract
Panoptic 3D reconstruction from a monocular video is a fundamental perceptual task in robotic scene understanding. However, existing efforts suffer from inefficiency in terms of inference speed and accuracy, limiting their practical applicability. We present EPRecon, an efficient real-time panoptic 3D reconstruction framework. Current volumetric-based reconstruction methods usually utilize multi-view depth map fusion to obtain scene depth priors, which is time-consuming and poses challenges to real-time scene reconstruction. To address this issue, we propose a lightweight module to directly estimate scene depth priors in a 3D volume for reconstruction quality improvement by generating occupancy probabilities of all voxels. In addition, compared with existing panoptic segmentation methods, EPRecon extracts panoptic features from both voxel features and corresponding image features, obtaining more detailed and comprehensive instance-level semantic information and achieving more accurate segmentation results. Experimental results on the ScanNetV2 dataset demonstrate the superiority of EPRecon over current state-of-the-art methods in terms of both panoptic 3D reconstruction quality and real-time inference. Code is available at https://github.com/zhen6618/EPRecon.
@@ I INTRODUCTION
Recently, visual-based panoptic 3D reconstruction [1] has gradually attracted attention due to its rich scene understanding information and wide application scenarios, such as robotic mapping and manipulation [2], and augmented reality (AR) [3]. Since camera poses can be accurately obtained by systems such as visual odometry [4], robot kinematics [5], or Simultaneous Localization and Mapping (SLAM) [6, 7], the main challenge in panoptic 3D reconstruction from a monocular video is how to obtain dense, high-quality scene surface depth from continuous RGB sequences and perform panoptic segmentation on the reconstructed scene.
@@
Methods that fuse multi-view depth maps obtained from similarity matching along epipolar lines, such as MVSNet series [8, 9, 10, 11, 12], suffer from depth estimation inconsistencies and redundant computations. To address this issue, volumetric-based methods [13, 14, 15, 16, 1] directly predict the explicit surface occupancy probabilities or implicit truncated signed distance function (TSDF) values to reconstruct scene surfaces in a 3D volume. To improve reconstruction quality, some works [17, 18, 1, 19] first use scene depth estimation models such as multi-view depth map fusion models [20] for a preliminary estimate of scene depth priors, which can eliminate the majority of non-surface voxels (over 90%), and then perform depth-guided reconstruction. Although these methods effectively improve reconstruction performance, the scene depth estimation networks used in existing methods are time-consuming, posing challenges to real-time scene reconstruction.
@@
To simultaneously obtain panoptic segmentation information and depth information from image sequences, recent studies [1, 21] first predict occupancy probabilities of voxels, and then conduct 3D panoptic segmentation on predicted reconstructed voxels. Although these methods have shown certain efficacy, their panoptic features are only extracted from occupied voxel features or corresponding image features, resulting in limited instance-level semantic information. This is because voxel features need to generate rich geometric features for depth predictions and usually emphasize detailed semantic information. Moreover, after multi-scale feature extraction and aggregation, image features generally emphasize comprehensive semantic information.
@@
In this paper, we propose EPRecon, an efficient real-time panoptic 3D reconstruction framework. We first estimate scene depth priors in a 3D volume to eliminate most of non-surface voxels, and then perform depth-guided panoptic 3D reconstruction. Unlike previous works [17, 18, 1, 19] which directly use time-consuming depth estimation models to obtain 2D depth priors and then perform depth fusion, EPRecon introduces a lightweight module to directly estimate scene depth priors in a 3D volume by generating occupancy probabilities of all voxels. Specifically, EPRecon back-projects multi-view image features into the volume and extracts feature similarities between views. These features are further aggregated and converted into occupancy probabilities of voxels. The module makes EPRecon focus more on voxel feature extraction in regions near scene surfaces and improves the entire panoptic reconstruction quality.
@@
  • We introduce a lightweight module to directly estimate scene depth priors in a 3D volume, which is significantly faster than multi-view depth map fusion-based methods and improves panoptic reconstruction quality.
  • Compared with current panoptic segmentation methods, we extract panoptic features from both voxel features and corresponding image features, obtaining more detailed and comprehensive panoptic segmentation information and achieving better segmentation performance.
@@ II Related Work
3D reconstruction from RGB image sequences has a long research history, and many classic works have emerged, such as MVSNet [8] and ORB-SLAM3 [6]. Multi-view stereo reconstruction methods represented by MVSNet series [8, 9, 10, 11, 12] estimate depth map for each view and then perform depth map fusion. These methods greatly improve the accuracy of 3D reconstruction, however, multiple depth estimations from different views suffer from depth estimation inconsistencies and redundant computations, usually generating artifacts and repeated surfaces. Volumetric-based approaches [13, 14, 15, 23, 24, 25, 26, 27] directly predict surface occupancy probabilities or TSDF values in a 3D volume. An early work is Atlas [13], which directly regressed a TSDF volume from a set of posed RGB images. NeuralRecon [14] reconstructed local surfaces represented by sparse TSDF volumes for each video fragment sequentially, and used GRUs to fuse TSDF volumes. To achieve more detailed reconstruction, recent studies [17, 18, 1, 19] first estimate scene depth priors and then perform depth-guided scene reconstruction. DG-Recon [17] utilized multi-view depth priors to eliminate many irrelevant non-surface voxels and guide surface reconstruction. FineRecon [18] combined depth priors and global image features for surface occupancy predictions. PanoRecon [1] performed panoptic reconstruction with the guidance of MVS depth estimation. Although a preliminary estimate of the scene depth can recover more accurate surfaces, the depth prior prediction modules in existing methods are time-consuming, posing challenges to real-time panoptic reconstruction. To address this issue, we propose a lightweight module to directly estimate scene depth priors in the 3D volume.
@@ IV-C Ablation Studies
In Table V, we conduct ablation studies to verify the effectiveness of the components in EPRecon. The qualitative comparison and analysis are shown in Fig. 3. The proposed depth prior estimation module effectively improves the quality of 3D geometric reconstruction and 3D panoptic segmentation. Since we directly estimate scene depth priors in a 3D volume, our proposed module is more lightweight (10.8$\times$ faster) than current depth map fusion-based methods used for depth prior estimation (see Table IV). In addition, we compare with extracting panoptic features only from voxel features or corresponding image features. Our method captures both types of features, having more detailed and comprehensive instance-level semantic information, which leads to higher 3D panoptic segmentation performance.
TABLE IV: Runtime analysis of panoptic 3D reconstruction. PanoRecon uses multi-view depth map fusion for depth prior estimation
@@
EPRecon 40 ms (10.8$\times$) 287 ms (2.4$\times$) 31.36 KFPS